the following investments you and your friend has provided their expectations for the markets for next year.

State of

Economy

Probability of

State of Economy

Stock

A

Stock

B

TSX

Boom

.30

30%

-9%

18%

Normal

.40

16%

12%

10%

Recession

.30

-10%

20%

-10%

              Remember to show all of your work - make clear and understandable and CORRECT

  1. Calculate the expected return for stock A.

  1. Calculate the expected return for stock B.

  

  1. Calculate the expected return for the TSX.
  1. Calculate the risk for stock A.
Answer #1

Expected Return = PB x RB + PN x RN + PR x RR

Where,

PB = Probability of Boom Economy

RB = Expected Return in a Boom Economy

PN = Probability of Normal Economy

RN = Expected Return in a Normal Economy

PR = Probability of Recession Economy

RR = Expected Return in a Recession Economy

Using the above formula,

Expected return for Stock A = EA = (0.3 x 30%) + (0.4 x 16%) + (0.3 x -10%) = 12.40%

Expected return for Stock B = (0.3 x -9%) + (0.4 x 12%) + (0.3 x 20%) = 8.10%

Expected return for the TSX = (0.3 x 18%) + (0.4 x 10%) + (0.3 x -10%) = 6.40%

In order to calculate the risk for Stock A, we need to find the standard deviation for Stock A which is calculated using the following formula:

Standard Deviation (Risk) = Square Root of [ PB x (RB - EA)2 + PN x (RN - EA)2 + PR x (RR - EA)2 ]

= Square Root of [ 0.3 x (30% - 12.4%)2 + 0.4 x (16% - 12.4%)2 + 0.3 x (-10% - 12.4%)2 ]

So, Risk for Stock A = 15.77%